# Improving the fact-checking performance of language models by relying on their entailment ability

### v1

@@ D.3 Additional observations from TBE-3:
  • We observed that XLNet fine-tuned with Llama entailed justifications gave the highest macro-recall (0.54) for the LIAR-RAW dataset. Many models, such as RoBERTa fine-tuned with Mistral ($MR:\textbf{0.47}$, $\sim\textbf{27.02}\%\uparrow$), Llama ($MR:\textbf{0.53}$, $\sim\textbf{43.24}\%\uparrow$), Gemma ($MR:\textbf{0.50}$, $\sim\textbf{35.14}\%\uparrow$), Qwen ($MR:\textbf{0.47}$, $\sim\textbf{27.02}\%\uparrow$), and Falcon ($MR:\textbf{0.43}$, $\sim\textbf{16.22}\%\uparrow$) entailed justification, XLNet fine-tuned with Mistral ($MR:\textbf{0.47}$, $\sim\textbf{27.02}\%\uparrow$), Llama ($MR:\textbf{0.54}$, $\sim\textbf{45.95}\%\uparrow$), Gemma ($MR:\textbf{0.43}$, $\sim\textbf{16.22}\%\uparrow$), Qwen ($MR:\textbf{0.48}$, $\sim\textbf{29.73}\%\uparrow$), and Falcon ($MR:\textbf{0.44}$, $\sim\textbf{18.92}\%\uparrow$) entailed justifications, Llama trained (with Llama entailed justifications) with LoRA+ adapter ($MR:\textbf{0.50}$, $\sim\textbf{35.14}\%\uparrow$) and Falcon trained (with Falcon entailed justifications) with LoRA+ adapter ($MR:\textbf{0.39}$, $\sim\textbf{5.41}\%\uparrow$) surpassed the best reported macro-recall score of the baselines ($MR:\textbf{0.37}$). Similarly, for the RAW-FC dataset, we observed that many models, such as such as RoBERTa fine-tuned with Mistral ($MR:\textbf{0.82}$, $\sim\textbf{34.43}\%\uparrow$), Llama ($MR:\textbf{0.88}$, $\sim\textbf{44.26}\%\uparrow$), Qwen ($MR:\textbf{0.71}$, $\sim\textbf{16.39}\%\uparrow$), and Falcon ($MR:\textbf{0.64}$, $\sim\textbf{4.92}\%\uparrow$) entailed justifications, XLNet fine-tuned with Mistral ($MR:\textbf{0.82}$, $\sim\textbf{34.43}\%\uparrow$), Llama ($MR:\textbf{0.88}$, $\sim\textbf{44.26}\%\uparrow$), Qwen ($MR:\textbf{0.70}$, $\sim\textbf{14.75}\%\uparrow$), and Falcon ($MR:\textbf{0.75}$, $\sim\textbf{22.95}\%\uparrow$) entailed justifications, Llama trained (with Llama entailed justifications) with LoRA ($MR:\textbf{0.62}$, $\sim\textbf{1.64}\%\uparrow$) and LoRA+ adapter ($MR:\textbf{0.82}$, $\sim\textbf{34.43}\%\uparrow$) surpassed the best reported macro-precision score of baselines ($MR:\textbf{0.61}$). RoBERTa and XLNet fine-tuned on Llama entailed justifications gave the highest overall performance ($MR:\textbf{0.88}$).
D.4 Additional observations from IBEs:
  • For LIAR-RAW dataset, we observed that Qwen gave the highest macro-precision (0.27) in IBE-1. While in IBE-2, Llama achieved the highest macro-precision ($MP:\textbf{0.27}$), Mistral gave the highest macro-precision in IBE-3 ($MP:\textbf{0.27}$) and IBE-4 ($MP:\textbf{0.27}$). However, none of them could surpass the baselines. In contrast, for RAW-FC dataset, Qwen ($MP:\textbf{0.61}$), Llama ($MP:\textbf{0.62}$), and Falcon ($MP:\textbf{0.60}$) achieved highest macro-precision in IBE-1, IBE-2, and IBE-3 respectively. They are comparable to the highest macro-precision reported by the baseline ($MP:$ 0.62). Similarly, Qwen ($MP:\textbf{0.50}$) got the best macro-precision in IBE-4, which is far behind the baselines.
@@ D.4 Additional observations from IBEs:
  • For LIAR-RAW dataset, Mistral achieved the highest macro-recall (0.25) in IBE-1. While in IBE-2, Mistral, Llama and Qwen achieved the highest macro-recall ($MR:\textbf{0.23}$), for IBE-3 Mistral and Falcon ($MR:\textbf{0.23}$) gave the highest scores. None of them could surpass the baseline performances. Similarly, they achieved highest scores in IBE-4 ($MR:\textbf{0.17}$). In contrast, for RAW-FC dataset, Llama ($MR:\textbf{0.63}$, $\sim\textbf{3.29}\%\uparrow$) achieved highest performace in IBE-2, surpassing the highest baseline performance ($MR:$ 0.61). While Qwen attained the highest performance in IBE-1 ($MR:$ 0.58), surpassing the baselines HiSS ($MR:$ 0.54), FactLLaMa ($MR:$ 0.55) and RAFTS ($MR:$ 0.52), it is still behind the performance of L-Defense ($MR:$ 0.61). Similarly, Qwen got the highest performance in IBE-3 ($MR:$ 0.54), which surpassed the baseline RAFTS ($MR:$ 0.52), but it is still behind the performance of HiSS ($MR:$ 0.54), FactLLaMa ($MR:$ 0.55), and L-Defense ($MR:$ 0.61). In IBE-4, Mistral and Qwen achieved the highest macro-recall of 0.46, which is far behind all baselines.
D.5 Additional observations from comparing the performance of TBEs and IBEs:
@@
  • We compared the macro-precision scores of best-performing models across TBEs and IBEs and presented them in Figure 23 for illustration. For the LIAR-RAW dataset, none of the models from IBEs, TBE-1, or TBE-2 ($MP$: 0.27 - 0.44) could exceed the highest macro-precision reported by the baseline ($MP$: 0.47). However, some models in TBE-1 and TBE-2 achieved higher macro-precision (0.36 - 0.44, $\sim\textbf{10.00}\%\uparrow$) than the IBEs ($MP$: 0.27 - 0.40). Moreover, several TBE-3 models ($MP$: 0.48 - 0.55, $\sim\textbf{25.00}\%\uparrow$) outperformed all models in IBEs ($MP$: 0.27 - 0.40), TBE-1 ($MP$: 0.44), and TBE-2 ($MP$: 0.36). For the RAW-FC dataset, we observed similar comparative results. Here, the highest macro-precision reported by an IBE-2 (0.62) model was nearly the same as the best-performing baseline ($MP$: 0.62). But, none of the models from IBEs ($MP$: 0.50 - 0.62) were able to surpass the baseline ($MP$: 0.62). Similar to the LIAR-RAW dataset, here we found that (i) many TBE-1 and TBE-2 models ($MP$: 0.63 - 0.73, $\sim\textbf{17.74}\%\uparrow$) outperformed the IBE models (0.50-0.62), and (ii) many TBE-3 models ($MP$: 0.63 - 0.88, $\sim\textbf{20.55}\%\uparrow$) outperformed all of the models in IBEs (0.50-0.62), TBE-1 ($MP$: 0.64 - 0.70) and TBE-2 ($MP$: 0.63 - 0.73).
  • We have also compared the macro-recall scores of the best-performing models across TBEs and IBEs and presented them in Figure 24. In the LIAR-RAW dataset, none of the models from IBEs, TBE-1 and TBE-2 ($MR$: 0.17 - 0.32) could outperform the highest baseline macro-recall (0.37). However, some models in TBE-1 and TBE-2 got better macro-recall (0.30 - 0.32, $\sim\textbf{28.00}\%\uparrow$) compared to the models in IBEs (0.25). Further, several TBE-3 models ($MR$: 0.43 - 0.54, $\sim\textbf{68.75}\%\uparrow$) outperformed models in IBEs ($MR$: 0.17 - 0.25), TBE-1 ($MR$: 0.30), and TBE-2 ($MR$: 0.32) in terms of the highest reported macro-recall. In the RAW-FC dataset, we observed a similar trend. The highest macro-recall reported by a model in IBE-2 (0.63, $\sim\textbf{3.28}\%\uparrow$) surpassed all baselines (0.61) by a small margin, whereas macro-recall from other IBEs remained significantly low (0.46 - 0.58). Consistent with LIAR-RAW results, (i) various TBE-1 and TBE-2 models ($MR$: 0.62 - 0.71, $\sim\textbf{12.70}\%\uparrow$) outperformed IBEs ($MR$: 0.46 - 0.63), and (ii) many TBE-3 models ($MR$: 0.64 - 0.88, $\sim\textbf{23.94}\%\uparrow$) surpassed all models in IBEs ($MR$: 0.46 - 0.63), TBE-1 ($MR$: 0.62 - 0.66) and TBE-2 ($MR$: 0.63 - 0.71), showing a significant improvement in terms of the highest reported macro-recall.
  • The confusion metrics for best-performing TBE and IBE models for LIAR-RAW and RAW-FC datasets are illustrated in Figure 37 and Figure 38, respectively. For LIAR-RAW dataset, TBE-3 recorded the highest number of correct predictions (CP) for the labels “true (CP : 186)”, “mostly-true (CP : 135)”, “false (CP : 195)”, “pants-fire (CP : 41)”. For “half-true (CP : 105)” IBE-4 achieved the most correct predictions, while TBE-2 performed best for “barely-true (CP : 115)”. Similarly, for RAW-FC dataset, TBE-3 recorded the highest number of correct predictions for the labels “true (CP : 60)”, “half (CP : 62)”, “false (CP : 55)”, surpassing other methods in correctly classifying these labels. These results showed that TBE-3 is effectively handling the label-wise distinction in the LIAR-RAW dataset. Here, CP denotes the number of correct predictions.
D.6 Detailed observations from evaluation of explanations:
In this section, we have reported our observations from the evaluation of model explanations. The details of evaluation metrics and approaches are reported in the section C.6.4. The results of lexical-overlapping and semantic-matching based evaluation are reported in Table 11. Similarly, the results of subjective evaluation are reported in Figure 26 and 27 (see Table 12 for raw values). Some of the key findings we got are,
  • Explanations generated by Falcon got highest $R_{1}$ score for both LIAR-RAW (0.23) and RAW-FC (0.40). It indicates that Falcon-generated explanations show maximum unigram overlap with the gold explanations provided in the datasets. Similarly, explanations generated by Mistral got the highest $R_{L}$ score for both LIAR-RAW (0.14) and RAW-FC (0.18). While Falcon too got the highest $R_{L}$ for LAIR-RAW (0.14), it fell behind for RAW-FC (0.17) by a small margin. Similar to Falcon, Llama also showcased a competitive $R_{L}$ score for RAW-FC (0.17). It indicates that explanations generated by these VLLMs show a maximum overlap of longest common subsequences with the gold explanations. Interestingly, we see a small deviation for $R_{2}$ scores. While explanations generated by Mistral, Llama and Falcon scored high $R_{2}$ for LIAR-RAW (0.06- 0.07) dataset, Gemma scored highest (0.20) for RAW-FC dataset. It means their explanations show maximum bigram overlap with the gold explanations. Except for the Gemma-generated explanations(BERT score: 0.24) for RAW-FC, BERT-scores were consistently low for all VLLMs and datasets (0.02-0.08). Similarly, BLEU scores were also low for both LIAR-RAW (0.02 - 0.03) and RAW-FC (0.02 - 0.07) datasets.
  • As evaluating VLLMs, four out of five models i.e. Mistral, Qwen, Llama and Falcon found that Llama-generated explanations are better in four out of five evaluating dimensions i.e. (informativeness: 4.73 - 4.78, readability: 4.51 - 4.91, objectivity: 4.17 - 4.60 and Logicality: 4.26 - 4.68) for the LIAR-RAW dataset. However, the majority (Llama, Gemma, Qwen and Falcon) of them found that Falcon-generated explanations are more accurate (3.55 - 3.96). Similarly, three out of five evaluating VLLMs, i.e. Mistral, Llama and Qwen, found that Llama-generated explanations are better for all five dimensions (informativeness: 4.48 - 4.92, accuracy: 4.10 - 4.14, readability: 4.43 - 4.82, objectivity: 4.28 - 4.47 and Logicality: 4.33 - 4.52). This observation seems to correlate with veracity prediction results, as RoBERTa and XLNet gave the high macro-precision, macro-recall and macro-F1 when trained with Llama-generated explanations in TBE-3.
  • We see some deviating trends as well. For the LIAR-RAW dataset, while most of the evaluating VLLMS gave high scores to Llama and Falcon-generated explanations, Gemma gave high scores to Mistral-generated explanations for three (informativeness: 3.70, objectivity: 3.91 and Logicality: 3.99) out of five dimensions. Similarly, for RAW-FC, we see that evaluating VLLMs Gemma and Falcon gave high scores to Mistral and Gemma-generated explanations. Figure 26 and Figure 27 indicate another interesting phenomenon. Gemma, as an evaluating VLLM, gave low scores compared to other evaluating VLLMs. We see a similar trend for Falcon as well in the RAW-FC dataset.
@@ D.6 Detailed observations from evaluation of explanations:
$R_{1}$ $R_{2}$ $R_{L}$ $BLEU$ $BERT$ $R_{1}$ $R_{2}$ $R_{L}$ $BLEU$ $BERT$
@@
Appendix E Detailed observations from ablation study:
Method LIAR-RAW RAW-FC
MP MR MF1 MP MR MF1
TBE-3 0.55 0.54 0.54 0.88 0.88 0.88
     -w/o Supp. just. 0.38 0.38 0.37 0.77 0.78 0.77
(±0.06) (±0.05) (±0.05) (±0.02) (±0.02) (±0.02)
     -w/o Ref. just. 0.49 0.50 0.49 0.80 0.80 0.80
(±0.01) (±0.02) (±0.02) (±0.02) (±0.02) (±0.02)
     -w/o Both just. 0.26 0.26 0.24 0.46 0.46 0.46
(±0.04) (±0.03) (±0.03) (±0.03) (±0.03) (±0.03)
In this section, we reported our observations from the ablation study. We took the best-performing models (XLNet-LLlama for LIAR-RAW and RoBERTa-LLlama for RAW-FC) for both LIAR-RAW and RAW-FC and removed individual justification (supporting and refuting) components to gauge its impact. Particularly, we trained the best-performing model frameworks by (i) removing supporting justification, (ii) removing refuting justification, and (iii) removing both. Results obtained for all training scenarios are reported in Table 13. Apart from that, we also segmented the test set based on the number of pieces of evidence each sample has, and calculated their performances. Particularly for LIAR-RAW, and RAW-FC, we divided the test set into six and five segments. The segment detail and their performance scores are reported in Table 14, and Table 15 respectively. We observed the following,
  • Upon removing the supporting justifications from training input (see (‘w/o Supp. just.’ in Table 13), we found that macro-precision, macro-recall and macro-F1 scores dropped by 30.90%, 29.63% and 31.48% respectively for the LIAR-RAW dataset. A similar trend can also be seen for the RAW-FC dataset, where macro-precision, macro-recall and macro-F1 scores dropped by 12.50%, 11.36% and 12.50% respectively. However, the performance drop was not as harsh as we saw for the LIAR-RAW dataset.
  • When we removed the refuting justifications from the training input (see (‘w/o ref. just.’ in Table 13), we saw a performance drop as well. We found that macro-precision, macro-recall and macro-F1 scores dropped by 10.90%, 7.41% and 9.26% respectively for the LIAR-RAW dataset. Similarly, macro-precision, macro-recall and macro-F1 scores dropped by 9.09%, 9.09% and 9.09% respectively for the RAW-FC dataset as well. One interesting observation is that removing the supporting justifications made a more adversarial impact than removing refuting justifications.
  • Finally, when we removed both supporting and refuting justifications from the training input, the models performed worst. The macro-precision, macro-recall and macro-F1 scores dropped by 52.73%, 51.85% and 55.56% respectively for the LIAR-RAW dataset. We observed a similar trend for RAW-FC as well. Here macro-precision, macro-recall and macro-F1 scores dropped by 47.73%, 47.73% and 47.73% respectively.
  • While investigating the behaviour with varying number of evidences, we observed the following. For LIAR-RAW, performance of veracity predictor peaked with ‘6–20’ evidences (MP: 0.62, MR: 0.60, MF1: 0.61). It showed significant sensitivity to increasing number of evidences with macro-F1 dropping $\textbf{29.50}\%$ from ‘6–20’ evidences (MF1: 0.61) to ‘>50’ evidences (MF1: 0.43). However, for RAW-FC, the highest were observed with ‘11–20’ evidences (MP: 0.93, MR: 0.95, MF1: 0.94). It showed more robustness with macro-F1 decreasing slightly $\textbf{7.45}\%$ from ‘11–20’ evidences (MF1: 0.94) to ‘>50’ evidences (MF1: 0.87).
@@ Appendix E Detailed observations from ablation study:
Appendix F Linguistic insights from model explanations:
In this section, we have reported how our best-performing model (XLNet-LLlama for LIAR-RAW and RoBERTa-LLlama for RAW-FC) is giving attention to different words and phrases while predicting veracity. We presented some of the samples, their gold labels, predicted labels, support justifications, and refute justifications in Figure 28, Figure 29, Figure 30, Figure 31, Figure 32, Figure 33, Figure 34, Figure 35, Figure 36. To visualise the attention, we highlighted the top $25\%$ words which received higher attention scores from the respective LLM in TBE-3. Here, the higher intensity of blue colour indicates a higher attention score. Some of the observations are,
  • In Figure 28 where the gold and predicted label is ‘true’, we observed high attention scores on the words like ‘supported by multiple sources’, ‘evidence’, ‘supports’ in the support justification part. However, attention scores are more scattered in refute justification. Also, in contrast to supporting justification, refuting justification doesn’t consist of any phrases which refutes the claim with authority.
  • In Figure 29 where the label is ‘mostly-true’, attention scores indicates that the LLM relied heavily on quantitative results which can be seen by attention on percentage scores in the supportive justification. But in refutive justification, LLM focused on counter-evidence but with less intensity of attention which matches the "mostly-true" stance.
  • In Figure 30 where the label is ‘half-true’, we observed the LLM could attend to some similar keywords in both support and refute justification. For example, ‘Gov’, ‘education’, etc. are prominent ones. Also, we observed more attention towards phrases like ‘appears that the claim is accurate’ in support justification and ‘inconsistencies undermine the validity of the claim’ in refute justification. It justifies the predicted label as half-true.
  • In Figure 31 where the label is ‘barely-true’, the VLLM generated justification focussing on references to the Obamacare and how many health plans were canceled nationwide. It noted Florida may have been heavily affected because of its size. But it also noticed that the exact number (300,000) is not confirmed. Subsequently, the LLM also gave more attention to keywords like ‘Obamacare’ due to which the LLM may have understood the claim might be true. Similarly, under the refute justification also, VLLM confidently contradicted the claim by saying that ‘customers were not immediately dropped’. As a result, the LLM extends more attention to ‘false’ and ‘coverage’, marking the reasons to doubt the claim. Thus, the LLM figured out that the claim is not entirely false, but also not fully accurate and labeled it as ‘barely-true’.
  • In Figure 32 where the label is ‘false’, the LLM assigned higher attention scores to economic indicators, due to absence of reasons to support the claim. Whereas the refuting justification consisted of statements disproving the claim and LLM showed more attention toward terms like ‘layoffs’, ‘assertion’, and ‘deaths’ which provides counter-evidence to refute the claim.
  • In Figure 33 where the label is ‘pants-fire’, the support justification is not useful as it is filled with repetition and irrelevant chat-bot fillers like ‘anything else I can help you with’, ‘let me know I can assist you further’, etc. Thus, we observed more attention scores on non-useful stopwords. While in the refuting justification the VLLM argued with authority that the claim is misleading and lacks merit. We observed high attention scores on the phrase ‘the claim is unfounded and lacks merit’ which aligns with the predicted label ‘pants-fire’.
  • In Figure 34 where the label is ‘true’, the support justification generated by VLLM emphasized on evidences like Fintan O’s (well-known columnist) writing and reputation of ‘Irish Times’. The LLM also assigned higher attention scores to these details which supports the claim with authority. However, refute justification seems to be weaker in disproving the claim due to lack of refuting evidence.
  • In Figure 35 where the label is ‘half’ (denoting half-true), the LLM put high attention on keywords like ‘evidence’, ‘statement is true’, ‘credibility’ and ‘support’ around factual components justifying truthfulness in the support justification. While in the refute justification, we observed higher attention scores on keywords like ‘false’, and phrases like ‘disputes the claim’ and ’In conclusion, the evidence suggests that the claim is false’. That is how the LLM evaluated the conflicting narratives and reached to a ‘half’ label.
  • In Figure 36 where the label is ‘false’, VLLM did not generate any factual justification for supporting the claim. Subsequently, the LLM could not find any solid fact or figure to put more attention on it. On the other hand, in the refute justification, VLLM justified with facts like event being from 2011, not during COVID. Also, the horses were rehomed well before the claim was made. Thus, the LLM could extend higher attention scores to phrases like ‘claim is not supported by facts’ and ‘outdated story’.

### v2

@@ D.1 Additional observations for macro precision in veracity prediction:
  • Several models within TBE-3 for LIAR-RAW demonstrated notable improvements over the best baseline macro-precision score ($MP$: 0.47). Specifically, (i) RoBERTa fine-tuned with Mistral ($\sim\textbf{2.12}\%\uparrow$), Llama ($\sim\textbf{12.76}\%\uparrow$), Gemma ($\sim\textbf{4.25}\%\uparrow$), Qwen ($\sim\textbf{2.12}\%\uparrow$), and Falcon ($\sim\textbf{4.25}\%\uparrow$) based entailed justifications, and (ii) XLNet fine-tuned with Mistral ($\sim\textbf{4.25}\%\uparrow$), Llama ($\sim\textbf{17.02}\%\uparrow$), and Qwen ($\sim\textbf{6.38}\%\uparrow$) based entailed justifications, and (iii) Llama trained with Llama justifications and LoRA+ adapter ($\sim\textbf{4.25}\%\uparrow$). Similarly, for RAW-FC, we observed that (i) RoBERTa fine-tuned with Mistral ($\sim\textbf{33.87}\%\uparrow$), Llama ($\sim\textbf{41.93}\%\uparrow$), Qwen ($\sim\textbf{17.74}\%\uparrow$), and Falcon ($\sim\textbf{4.83}\%\uparrow$) based entailed justifications, (ii) XLNet fine-tuned with Mistral ($\sim\textbf{33.87}\%\uparrow$), Llama ($\sim\textbf{41.93}\%\uparrow$), Qwen ($\sim\textbf{12.90}\%\uparrow$), and Falcon ($\sim\textbf{22.58}\%\uparrow$) based entailed justifications, and (iii) Llama trained (with Llama justifications) with LoRA+ adapter ($\sim\textbf{35.48}\%\uparrow$) outperformed the best reported macro-precision by the baselines ($MP$: 0.62) Models from TBE-3, exhibit the highest overall macro precision performance. Figure 23 illustrates the same.
D.2 Additional observations for macro recall in veracity prediction:
We reported the $MR$ of various models under TBE-1, TBE-2TBE-3, and IBEs in Table 11, Table 12, and Table 13, respectively. Figure 24 illustrates a comparison of the best-performing models across them. Our observations are as follows:
@@ D.2 Additional observations for macro recall in veracity prediction:
  • For LIAR-RAW, none of the macro-recall scores reported by IBE models could surpass the best baseline performance ($MR$: 0.37). In contrast, for RAW-FC, Llama achieved the highest performance in IBE-2 ($MR$: 0.63, $\sim\textbf{3.28}\%\uparrow$), surpassing the highest baseline performance ($MR$: 0.61).
  • In TBE-1, for LIAR-RAW, the highest reported $MR$ (0.30) does not surpass that of any baseline. However, for RAW-FC, we found that many models, such as (i) Mistral ($\sim\textbf{6.56}\%\uparrow$) and Qwen ($\sim\textbf{8.20}\%\uparrow$) trained with LoRA, and (ii) Llama trained with both LoRA ($\sim\textbf{4.92}\%\uparrow$) and LoRA+ ($\sim\textbf{4.92}\%\uparrow$), outperform the best baseline performance ($MR$: 0.61).
  • In TBE-2, for LIAR-RAW, no model surpasses the best $MR$ reported by the baselines. However, for the RAW-FC dataset, we observed that many models, such as (i) XLNet fine-tuned on Llama understandings ($\sim\textbf{3.27}\%\uparrow$), and (ii) Llama trained with Llama understandings and LoRA+ ($\sim\textbf{16.39}\%\uparrow$), outperformed the best reported baseline MRMR (MRMR: 0.61).
  • Several models in TBE-3 for LIAR-RAW, such as (i) RoBERTa fine-tuned with Llama ($\sim\textbf{43.24}\%\uparrow$), Gemma ($\sim\textbf{35.14}\%\uparrow$), and Falcon ($\sim\textbf{16.22}\%\uparrow$) based entailed justifications, (ii) XLNet fine-tuned with Llama ($\sim\textbf{45.95}\%\uparrow$), Qwen ($\sim\textbf{29.73}\%\uparrow$), and Falcon ($\sim\textbf{18.92}\%\uparrow$) based entailed justifications, and (iii) Llama trained with Llama justifications and LoRA+ adapter ($\sim\textbf{35.14}\%\uparrow$), surpassed the best reported macro-recall score of baselines ($MR$: 0.37). Similarly, for RAW-FC, we observed that (i) RoBERTa fine-tuned with Mistral ($\sim\textbf{34.43}\%\uparrow$), Llama ($\sim\textbf{44.26}\%\uparrow$), and Qwen ($\sim\textbf{16.39}\%\uparrow$) based entailed justifications, (ii) XLNet fine-tuned with Mistral ($\sim\textbf{34.43}\%\uparrow$), Llama ($\sim\textbf{44.26}\%\uparrow$), Qwen ($\sim\textbf{14.75}\%\uparrow$), and Falcon ($\sim\textbf{22.95}\%\uparrow$) based entailed justifications, and (iii) Llama trained (with Llama justifications) with LoRA+ adapter ($\sim\textbf{34.43}\%\uparrow$) outperformed the best reported macro-recall score by the baselines ($MR$: 0.61). Models from TBE-3, exhibit the highest overall macro recall performance. Figure 24 illustrates the same.
[…]
TBE-1 TBE-1
Method ($\downarrow$) MP MR MP MR
LoRA
     -Mistral 0.44 0.29 0.69 0.65
(±0.02) (±0.01) (±0.01) (±0.00)
     -Llama 0.34 0.30 0.68 0.64
(±0.01) (±0.02) (±0.01) (±0.02)
     -Gemma 0.27 0.25 0.60 0.58
(±0.02) (±0.03) (±0.05) (±0.04)
     -Qwen 0.37 0.29 0.67 0.66
(±0.02) (±0.02) (±0.03) (±0.03)
     -Falcon 0.39 0.28 0.59 0.58
(±0.01) (±0.01) (±0.03) (±0.05)
LoRA+
     -Mistral 0.40 0.27 0.53 0.56
(±0.03) (±0.02) (±0.05) (±0.03)
     -Llama 0.34 0.29 0.67 0.64
(±0.01) (±0.01) (±0.01) (±0.01)
     -Gemma 0.27 0.23 0.61 0.57
(±0.02) (±0.02) (±0.03) (±0.02)
     -Qwen 0.36 0.29 0.70 0.65
(±0.01) (±0.01) (±0.02) (±0.02)
     -Falcon 0.37 0.30 0.64 0.62
(±0.02) (±0.01) (±0.03) (±0.03)
Table 11: Performance of LoRA and LoRA+ models on TBE-1 (LIAR-RAW and RAW-FC datasets), showing MP(macro precision) and MR(macro recall).
[…]
Method ($\downarrow$) TBE-2 TBE-3 TBE-2 TBE-3
MP MR MP MR MP MR MP MR
Fine-tuning
     -RoBERTa-LMistral 0.28 0.26 0.48 0.47 0.51 0.50 0.83 0.82
(±0.01) (±0.01) (±0.01) (±0.00) (±0.01) (±0.01) (±0.00) (±0.01)
     -RoBERTa-LLlama 0.27 0.28 0.53 0.53 0.50 0.50 0.88 0.88
(±0.02) (±0.01) (±0.01) (±0.01) (±0.01) (±0.01) (±0.01) (±0.01)
     -RoBERTa-LGemma 0.28 0.27 0.49 0.50 0.51 0.51 0.50 0.49
(±0.02) (±0.01) (±0.01) (±0.02) (±0.04) (±0.03) (±0.02) (±0.01)
     -RoBERTa-LQwen 0.30 0.28 0.48 0.47 0.51 0.49 0.73 0.71
(±0.01) (±0.02) (±0.01) (±0.02) (±0.01) (±0.01) (±0.02) (±0.02)
     -RoBERTa-LFalcon 0.29 0.28 0.49 0.43 0.50 0.48 0.65 0.64
(±0.02) (±0.02) (±0.01) (±0.02) (±0.00) (±0.01) (±0.02) (±0.02)
     -XLNet-LMistral 0.29 0.28 0.49 0.47 0.62 0.61 0.83 0.82
(±0.02) (±0.01) (±0.00) (±0.01) (±0.02) (±0.01) (±0.01) (±0.01)
     -XLNet-LLlama 0.31 0.29 0.55 0.54 0.63 0.63 0.88 0.88
(±0.02) (±0.01) (±0.01) (±0.02) (±0.03) (±0.03) (±0.01) (±0.01)
     -XLNet-LGemma 0.26 0.26 0.47 0.43 0.51 0.50 0.47 0.47
(±0.02) (±0.02) (±0.04) (±0.05) (±0.01) (±0.01) (±0.02) (±0.02)
     -XLNet-LQwen 0.29 0.29 0.50 0.48 0.60 0.58 0.70 0.70
(±0.02) (±0.02) (±0.01) (±0.01) (±0.02) (±0.01) (±0.03) (±0.03)
     -XLNet-LFalcon 0.26 0.26 0.45 0.44 0.61 0.60 0.76 0.75
(±0.03) (±0.01) (±0.01) (±0.02) (±0.02) (±0.01) (±0.01) (±0.01)
LoRA
     -Mistral 0.36 0.32 0.35 0.34 0.61 0.60 0.69 0.59
(±0.01) (±0.01) (±0.04) (±0.03) (±0.04) (±0.04) (±0.05) (±0.05)
     -Llama 0.31 0.27 0.36 0.32 0.48 0.48 0.63 0.62
(±0.04) (±0.03) (±0.03) (±0.03) (±0.06) (±0.03) (±0.03) (±0.03)
     -Gemma 0.21 0.22 0.26 0.23 0.47 0.45 0.36 0.34
(±0.01) (±0.01) (±0.03) (±0.03) (±0.03) (±0.03) (±0.02) (±0.01)
     -Qwen 0.32 0.27 0.40 0.36 0.56 0.54 0.48 0.46
(±0.09) (±0.02) (±0.07) (±0.07) (±0.07) (±0.00) (±0.03) (±0.03)
     -Falcon 0.28 0.26 0.27 0.24 0.56 0.57 0.43 0.40
(±0.01) (±0.00) (±0.03) (±0.03) (±0.03) (±0.01) (±0.01) (±0.03)
LoRA+
     -Mistral 0.33 0.30 0.30 0.32 0.66 0.60 0.52 0.48
(±0.01) (±0.01) (±0.07) (±0.02) (±0.05) (±0.05) (±0.06) (±0.04)
     -Llama 0.27 0.24 0.49 0.50 0.73 0.71 0.84 0.82
(±0.01) (±0.02) (±0.03) (±0.02) (±0.02) (±0.02) (±0.01) (±0.03)
     -Gemma 0.23 0.21 0.34 0.32 0.53 0.51 0.52 0.48
(±0.03) (±0.02) (±0.04) (±0.05) (±0.02) (±0.02) (±0.03) (±0.04)
     -Qwen 0.27 0.27 0.33 0.34 0.57 0.48 0.52 0.49
(±0.04) (±0.03) (±0.03) (±0.03) (±0.07) (±0.03) (±0.06) (±0.09)
     -Falcon 0.33 0.28 0.39 0.39 0.55 0.56 0.58 0.52
(±0.00) (±0.01) (±0.02) (±0.03) (±0.03) (±0.02) (±0.01) (±0.02)
Table 12: Performance of TBE-2 and TBE-3 for the claim veracity prediction using gold evidences. Green and Blue indicate best and second-best performance, respectively.
[…]
Method ($\downarrow$) IBE-1 IBE-2 IBE-3 IBE-4 IBE-1 IBE-2 IBE-3 IBE-4
MP MR MP MR MP MR MP MR MP MR MP MR MP MR MP MR
Prompting
     -Mistral 0.24 0.25 0.22 0.23 0.40 0.23 0.30 0.17 0.54 0.54 0.58 0.58 0.50 0.52 0.45 0.43
     -Llama 0.24 0.23 0.30 0.23 0.26 0.22 0.22 0.13 0.56 0.56 0.62 0.63 0.45 0.47 0.42 0.39
     -Gemma 0.24 0.21 0.16 0.16 0.27 0.19 0.09 0.16 0.40 0.40 0.41 0.38 0.45 0.41 0.27 0.31
     -Qwen 0.27 0.23 0.25 0.23 0.24 0.22 0.13 0.16 0.61 0.58 0.58 0.57 0.57 0.54 0.50 0.46
     -Falcon 0.24 0.23 0.22 0.22 0.24 0.23 0.16 0.17 0.56 0.57 0.60 0.59 0.60 0.52 0.40 0.38
Table 13: Performance of Prompting methods across IBE-1, IBE-2, IBE-3, and IBE-4 settings for LIAR-RAW and RAW-FC datasets, showing MP and MR.
@@
Method LIAR-RAW RAW-FC
MP MR MF1 MP MR MF1
TBE-3 0.55 0.54 0.54 0.88 0.88 0.88
     -w/o Supp. just. 0.38 0.38 0.37 0.77 0.78 0.77
(±0.06) (±0.05) (±0.05) (±0.02) (±0.02) (±0.02)
     -w/o Ref. just. 0.49 0.50 0.49 0.80 0.80 0.80
(±0.01) (±0.02) (±0.02) (±0.02) (±0.02) (±0.02)
     -w/o Both just. 0.26 0.26 0.24 0.46 0.46 0.46
(±0.04) (±0.03) (±0.03) (±0.03) (±0.03) (±0.03)
@@
$R_{1}$ $R_{2}$ $R_{L}$ BLEUBLEU BERTBERT $R_{1}$ $R_{2}$ $R_{L}$ BLEUBLEU BERTBERT